# arXiv:2603.00637v1[cond-mat.mtrl-sci] 28 Feb 2026

### v1

@@ II.3.2 Localized resolution of identity technique
For QSGWGW, similar to our implementation of RPA Shi et al. (2024, 2025) and $G^{0}W^{0}$ Ren et al. (2021a); Zhang et al. (2026), we employ an auxiliary basis set to represent the two-electron quantities and LRI to decompose the four-index tensors into three- and two-index ones. In this approach, a key step is to expand the products of NAOs in terms of auxiliary basis functions (ABFs):
@@
where $P_{\mu\mathbf{R}}((\mathbf{r}))$ denotes the ABFs which are required to be centered on atoms $I$ or $J$, on which the two NAOs are located. In Eq. 14, $C_{i(\mathbf{R}),j(\mathbf{R}^{\prime})}^{\mu(\mathbf{R})}$ are the LRI expansion coefficients, and the lattice vector $\mathbf{R}$ in parentheses specifies the unit cell where the NAO $i$ or ABF $\mu$ is localized in a periodic system.
@@ II.3.3 Space-time formalism of QSGWGW with NAO and LRI
$\Xi_{p\mathbf{k}}(\tau)=e^{-(\epsilon_{p\mathbf{k}}-\epsilon_{F})\tau}\left[\theta(\tau)(1-f_{p\mathbf{k}})-\theta(-\tau)f_{p\mathbf{k}}\right].$ (17)
Here $\theta(\tau)$ is the Heaviside step function, $f_{p\mathbf{k}}$ is the occupation number, and $\epsilon_{F}$ is the Fermi energy (chemical potential). For the sake of simplicity, only the spin-degenerate case with integer orbital occupations is considered. The extension to the spin polarized case is straightforward.
@@
$\chi_{\mu\nu}^{0}(\mathbf{R},\mathrm{i}\tau)=-\mathrm{i}\sum_{ijkl}\sum_{\mathbf{R}_{1},\mathbf{R}_{2}}C_{i(\mathbf{0}),k(\mathbf{R}_{1})}^{\mu(\mathbf{0})}C_{j(\mathbf{R}),l(\mathbf{R}+\mathbf{R}_{2})}^{\nu(\mathbf{R})}$
$\times\Big[G_{ij}^{0}(\mathbf{R},\mathrm{i}\tau)G_{lk}^{0}(\mathbf{R}_{1}-\mathbf{R}-\mathbf{R}_{2},-\mathrm{i}\tau)$
$\qquad+G_{ij}^{0}(\mathbf{R},-\mathrm{i}\tau)^{*}G_{lk}^{0}(\mathbf{R}_{1}-\mathbf{R}-\mathbf{R}_{2},\mathrm{i}\tau)^{*}\Big].$ (19)
@@
Now, we can calculate the correlation part of the self-energy in the NAO basis as follows:
$\Sigma_{pp^{\prime}}^{\mathrm{c}}(\mathbf{k},\mathrm{i}\tau)$ $=\sum_{ij}c_{p\mathbf{k}}^{i*}c_{p^{\prime}\mathbf{k}}^{j}\sum_{\mathbf{R}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{R}}\Sigma_{ij}^{\mathrm{c}}(\mathbf{R},\mathrm{i}\tau).$ (22)
Here, the correlation matrix elements $\Sigma^{\rm c}_{ij}(\mathbf{R},\mathrm{i}\tau)$ can be expressed in terms of LRI coefficients, the screened Coulomb interaction $W_{\mu\nu}^{c}(\mathbf{R},\mathrm{i}\tau)$, and the Green’s function as:
@@
Here, the $GW$ self-energy is decomposed as $\Sigma=\Sigma^{\rm x}+\Sigma^{\rm c}$. The correlation part uses $W_{\mu\nu}^{c}\equiv W_{\mu\nu}^{0}-V_{\mu\nu}$, whereas the exchange part is obtained from the same LRI machinery by using the bare Coulomb interaction $V_{\mu\nu}$ and is therefore omitted here; see Ref. Zhang et al. (2026) for implementation details.
@@ III.1 QSGWGW in LibRPA
Subsequently, the QSGWGW iteration proceeds analogously to a standard $G^{0}W^{0}$ calculation in LibRPA until the self-energy on the imaginary frequency axis, $\Sigma_{ij}(\mathbf{k},\mathrm{i}\omega)$, is obtained.
A Padé approximation is then employed to perform the analytic continuation from the imaginary frequency axis to the real frequency axis, yielding the self-energy at real frequencies, $\Sigma_{ij}(\mathbf{k},\omega_{\mathrm{real}})$. A detailed discussion on the stability of this analytic continuation is provided in Sec. III.2.
@@
To improve the stability of the self-consistency cycle for molecular systems, we employ Pulay’s DIIS mixing method Pulay (1980) when updating the QSGWGW Hamiltonian. For periodic systems, the iterations are generally stable in our benchmarks and converge without DIIS mixing. Our QSGWGW implementation typically requires 6–10 iterations to achieve convergence of eigenvalues within $10^{-3}$ eV for periodic systems, whereas molecular systems may still require tens to hundreds of iterations. Upon convergence, the final quasiparticle energies and orbitals are utilized to compute various electronic properties, such as band structures and topological invariants.
@@ III.2 Analytic continuation of self-energy
Figure 2: Basis representation dependence tests of analytic continuation in QSGWGW for “Mode A” and “Mode B” for Si (top) and MgO (bottom). The Si calculations employed a $6\times 6\times 6$ k-point mesh, while MgO employed an $8\times 8\times 8$ mesh. Both used a frequency grid of 16 points and the intermediate_gw basis sets. In the figure, “initial” denotes analytic continuation performed in the initial KS-DFT eigenbasis, while “updated” denotes analytic continuation performed in the updated QSGWGW eigenbasis from the previous iteration.
@@ III.3.1 $\mathbf{k}$-point and frequency grid convergence
First, we test the $\bf k$} convergence of the Si QSGWGW band gap with respect to Brillouin-zone sampling using uniform $n\times n\times n$ $\bf k$}-point meshes (n=6n=6–10), as shown in Fig. 3. Taking the $10\times 10\times 10$ mesh as the reference, the residual QSGWGW band-gap difference decreases from 32 meV at $6\times 6\times 6$ to 6.5 meV at $9\times 9\times 9$ (with $E_{g}^{\mathrm{QS}GW}=1.3621$ eV at $9\times 9\times 9$ and $E_{g}^{\mathrm{QS}GW}=1.3686$ eV at $10\times 10\times 10$, i.e., a last-step change of 6.46 meV); at the adopted $8\times 8\times 8$ mesh the residual relative to $10\times 10\times 10$ is 12.5 meV, but this level of k-point convergence is still not fully satisfactory.
In the present implementation, a likely reason for this residual k-point sensitivity is that the head correction term (i.e., the $\Gamma$-point correction to the dielectric matrix; see Ref. Zhang et al. (2026)) is updated only in the first QSGWGW iteration and is kept fixed in subsequent iterations. For Si at the adopted $8\times 8\times 8$ mesh, we estimate the associated residual uncertainty in $E_{g}$ to be on the order of 10–15 meV under the present protocol. We are continuing to improve this aspect by updating the treatment of the head correction term at each iteration.
@@ III.3.2 Basis sets
Here, intermediate_gw and really_tight denote the predefined FHI-aims species_defaults setting (defaults_2020), which explicitly specify (i) the orbital NAO basis (minimal basis plus additional confined radial functions, organized in tiers) and (ii) the associated numerical settings (integration grids and Hartree-potential accuracy) for each chemical element. In the defaults_2020 hierarchy, intermediate setting retains most numerical settings from tight but employs a more economical orbital basis, whereas tight setting corresponds to tier-based NAO orbital lists (e.g., ”tier 2” for H–Ne and ”tier 1” for most heavier elements), and really_tight setting keeps the same orbital basis as tight setting but uses strongly overconverged numerical parameters (including basis_dep_cutoff=0).FHI-aims team (2025)
In the corresponding RI-based $GW$ workflow, the auxiliary basis refers to the atom-centered auxiliary basis functions used to expand products of orbital basis functions in the Coulomb-operator representation.
For periodic $GW$ calculations, the corresponding intermediate_gw setting is identical to the standard intermediate setting except that it augments the ABFs used in the LRI (RI-LVL) representation of the Coulomb operator by adding extra for_aux functions (specifically, for_aux hydro 4 f 0.0); these for_aux functions are used only to generate the auxiliary basis and do not enter the KS orbital basis.FHI-aims team (2025); Ren et al. (2021b)
@@
Table 1: Experimental vertical ionization potentials (IP, in eV), together with our QSGWGW results and independent QSGWGW reference values from two GW100 external datasets (ADF STO extra and TM def2-TZVPP). For GWGW-based methods, IPs correspond to $\varepsilon_{\mathrm{H}}^{\mathrm{QP}}$, where $\varepsilon_{\mathrm{H}}^{\mathrm{QP}}$ is the converged HOMO quasiparticle energy. Experimental IPs are those compiled in Ref. van Setten et al. (2015) (Table 2 therein; values marked * are vertical, following that compilation). Each molecule employs the really_tight NAO basis tier and 32 frequency points. The external QSGWGW reference values are taken from the GW100 database 33 (see main text for details). Entries marked “—” indicate missing values.
Both settings correspond to predefined FHI-aims NAO tiers and have been shown to yield well-converged $G^{0}W^{0}$ quasiparticle gaps for NAO-based all-electron implementations on similar benchmark sets. We therefore adopt the same basis tiers for QSGWGW to provide a consistent and well-tested numerical baseline for the present benchmarks.
@@ IV.1 Molecular Benchmarks
To validate our QSGWGW implementation for molecular systems, we perform benchmark calculations on a selection of small molecules from the GW100 test set van Setten et al. (2015); Kaplan et al. (2016); Caruso et al. (2016); van Setten et al. (2013).
As an initial validation of the present implementation, we consider a representative subset of small molecules from the GW100 test set (Table 1). The selected molecules (Ag2, AlF3, As2, AsH3, BF, BH3, and C2H2) cover diverse bonding characters and chemical environments, including a metallic dimer (Ag2), an ionic fluoride (AlF3), $p$-block elemental/covalent species (As2, AsH3), a polar diatomic (BF), an electron-deficient hydride (BH3), and a small $\pi$-bonded hydrocarbon (C2H2). Together, they provide a compact yet nontrivial validation set for the present implementation.
@@
In Table 1, the GW100 reference datasets are labeled by the underlying code and basis setting. In particular, “ADF STO extra” refers to the QSGWGW HOMO dataset computed with the Amsterdam Density Functional (ADF) code using a Slater-type orbital (STO) basis in the larger “extra” basis sets Förster and Visscher (2021) as provided in the GW100 database (GW100 file qsGW_HOMO_ADF_STO_extra.json). Likewise, “TM def2-TZVPP” refers to the QSGWGW HOMO dataset computed with TURBOMOLE (TM) using the Gaussian def2-TZVPP basis sets Caruso et al. (2016) as provided in the GW100 database (GW100 file qsGW_HOMO_Tv6.0_def2-TZVPP.json).
@@ IV.2 Semiconductors and Insulators
wAlN wurtzite 3.110 4.980
wZnO wurtzite 3.250 5.207
@@
Table 2: Crystal structure prototypes and lattice constants (in Å) used for the periodic benchmark set in Table 3. For cubic systems, $c$ is not applicable.
Table 3 presents the band gaps of a series of semiconductors and insulators computed by our NAO-based QSGWGW implementation. We first prioritize validation against established implementations by comparing our QSGWGW gaps to literature benchmarks from independent codes (Table 3). In particular, we list QSGW reference values obtained with crystalline Gaussian basis sets Lei and Zhu (2022) and all-electron LAPW-based QSGW benchmark values Salas-Illanes et al. (2022). For additional context on the spread across widely used implementations, we also list representative self-consistent GW-family reference results obtained within a PMT (plane waves plus muffin-tin orbitals) hybrid-QSGW workflow Deguchi et al. (2016), a PAW-based fully self-consistent GW implementation Grumet et al. (2018b), and a plane-wave self-consistent GW scheme with efficient vertex corrections Chen and Pasquarello (2015).
To facilitate cross-code validation, we treat the Gaussian-QSGW Lei and Zhu (2022) and LAPW-QSGW Salas-Illanes et al. (2022) columns in Table 3 as our primary QSGWGW reference benchmarks where available. The remaining literature columns (PMT hQSGW Deguchi et al. (2016), PAW scGW Grumet et al. (2018b), and PW scGW+vertex Chen and Pasquarello (2015)) represent closely related self-consistent $GW$-family workflows and are included to make the spread across widely used approaches explicit; since they are not strictly at the same approximation level as standard QSGWGW, we use them as context rather than as one-to-one QSGWGW benchmarks. The superscript $\mathrm{zb}$ denotes that the corresponding ZnO value was reported for the zinc-blende structure in Ref. Grumet et al. (2018b) (as compiled in Ref. Lei and Zhu (2022)).
@@
Gaussian QSGW Lei and Zhu (2022) LAPW QSGW Salas-Illanes et al. (2022) PMT hQSGW Deguchi et al. (2016) PAW scGW Grumet et al. (2018b) PW scGW+vertex Chen and Pasquarello (2015) FP-LMTO QSGW Kotani et al. (2007)
@@
MPE (%) -45.64 -10.05 ++9.43
Table 3: Band gaps (in eV) of semiconductors and insulators. An $8\times 8\times 8$ k-point mesh was employed for most systems, with the exception of wAlN and wZnO ($8\times 8\times 5$). All calculations utilized a frequency grid of 16 points and the intermediate_gw basis sets (see Sec. III.3.2). Experimental band gaps are taken from Ref. Ren et al. (2021a) (and references therein), except for Ne which is taken from Ref. Schwentner et al. (1975). For BP, reported experimental values vary (2.2–2.5 eV); we use the midpoint 2.35 eV when computing the error metrics. Literature reference values are cited in the column headers (see main text for interpretation details). The Mean Absolute Relative Error (MARE) and Mean Percentage Error (MPE) with respect to experimental values are listed at the bottom for PBE, $G^{0}W^{0}$@PBE, and our QSGWGW results. The MPE indicates systematic underestimation (negative values) or overestimation (positive values). All values are reported to two decimal places.
@@
As illustrated in Fig. 5, the PBE functional systematically underestimates the band gaps for this benchmark set. The single-shot $G^{0}W^{0}$@PBE method captures most of the gap opening relative to PBE, while QSGWGW provides a further systematic correction on top of $G^{0}W^{0}$@PBE under the present protocol. In what follows, we validate our QSGWGW gaps against independent QSGWGW benchmarks from the literature, and list experimental gaps only as context.
@@
First, we assess cross-implementation consistency by comparing our QSGWGW gaps with literature QSGWGW benchmarks from independent implementations, focusing on Gaussian-QSGW Lei and Zhu (2022) and LAPW-QSGW Salas-Illanes et al. (2022) (Table 3). We define $\Delta E_{g}\equiv E_{g}^{\mathrm{(this\ work)}}-E_{g}^{\mathrm{(ref)}}$ in eV and report it for each system in Table 3. For the systems available in common, the deviations are typically within $\sim$0.3 eV, with the largest discrepancies for wZnO ($\Delta E_{g}=-0.92$ eV relative to Gaussian-QSGW) and MgO ($\Delta E_{g}=-0.83$ eV relative to LAPW-QSGW). Overall, the cross-code deviations are comparable to the spread seen among published periodic QSGWGW benchmarks and support the validity of the present NAO-based QSGWGW workflow at the level of established implementations.
Similarly, for wide-gap insulators like C (diamond) and BN, our results (6.41 eV and 7.29 eV) are consistent with the literature benchmarks listed in Table 3. This agreement suggests that our implementation, utilizing the space-time formalism and LRI technique within the NAO framework, captures the essential QSGWGW physics at a level comparable to established implementations used in published periodic benchmarks.
A notable deviation is observed for wurtzite ZnO (wZnO), a system known to be challenging for $GW$ methods. Our calculated QSGWGW band gap is 3.71 eV, which is lower than several established cross-code reference values reported for ZnO in the literature (Table 3). wZnO is notoriously difficult to converge in $GW$ calculations due to the strong hybridization between Zn $3d$ states and O $2p$ states, as well as the slow convergence with respect to high-energy unoccupied states.
Interestingly, while $G^{0}W^{0}$@PBE shows a large underestimation for wZnO (33.3% error), our QSGWGW result overestimates by only 3.1%, indicating that quasiparticle self-consistency improves the gap relative to $G^{0}W^{0}$@PBE for this material within the present numerical protocol. Previous studies have shown that, in the LAPW framework, adding high-energy local orbitals (HLOs) can significantly increase the $GW$ band gap of ZnO Jiang and Blaha (2016). The underestimation of the $G^{0}W^{0}$ band gap for ZnO within the standard NAO basis sets has been discussed in Ref. Ren et al. (2021b). It has been shown Ren et al. (2021b) that the issue can be mitigated by complementing NAOs with highly localized Slater-type orbitals (STOs), in analogy with the LAPW+HLOs scenario.
For the rare gas solid Ne, our QSGWGW result (22.84 eV) is in excellent agreement with the periodic QSGW benchmark value (22.57 eV) reported in Ref. Lei and Zhu (2022) and with the experimental value (21.7 eV). This confirms that our implementation reliably describes wide-gap systems with diverse bonding characteristics, from covalent semiconductors to weakly interacting rare gas solids.

### v2

@@ II.3.2 Localized resolution of identity technique
For QSGWGW, we employ an auxiliary basis set to represent the two-electron quantities and LRI to decompose the four-index tensors into three- and two-index ones, similar to the canonical $G^{0}W^{0}$ implementation for periodic systems in FHI-aims Ren et al. (2021) and our low-scaling RPA and $G^{0}W^{0}$ implementation in LibRPA Shi et al. (2024, 2025); Zhang et al. (2026). In this approach, a key step is to expand the products of NAOs in terms of auxiliary basis functions (ABFs):
@@
where $P_{\mu\mathbf{R}}(\mathbf{r})$ denotes the ABFs which are required to be centered on atoms $I$ or $J$, on which the two NAOs are located. In Eq. 14, $C_{i(\mathbf{R}),j(\mathbf{R}^{\prime})}^{\mu(\mathbf{R})}$ are the LRI expansion coefficients, and the lattice vector $\mathbf{R}$ in parentheses specifies the unit cell where the NAO $i$ or ABF $\mu$ is localized in a periodic system.
@@ II.3.3 Space-time formalism of QSGWGW with NAO and LRI
$\Xi_{p\mathbf{k}}(\tau)=e^{-(\epsilon_{p\mathbf{k}}-\epsilon_{\rm F})\tau}\left[\theta(\tau)(1-f_{p\mathbf{k}})-\theta(-\tau)f_{p\mathbf{k}}\right].$ (17)
Here $\theta(\tau)$ is the Heaviside step function, $f_{p\mathbf{k}}$ is the occupation number, and $\epsilon_{\rm F}$ is the Fermi energy (chemical potential). For the sake of simplicity, only the spin-degenerate case with integer orbital occupations is considered. The extension to the spin polarized case is straightforward.
@@
$\chi^{0}_{\mu\nu}(\mathbf{R},\mathrm{i}\tau)$ $=-\sum_{ij}\sum_{kl}\sum_{\mathbf{R}_{1},\mathbf{R}_{2}}C^{\mu(\mathbf{0})}_{i(\mathbf{0}),k(\mathbf{R}_{1})}\,C^{\nu(\mathbf{R})}_{j(\mathbf{R}),l(\mathbf{R}_{2})}$
$\quad\times\Big[G^{0}_{ij}(\mathbf{R},\mathrm{i}\tau)\,G^{0}_{lk}(\mathbf{R}_{1}-\mathbf{R}_{2},-\mathrm{i}\tau)$
$\qquad+G^{0}_{il}(\mathbf{R}_{2},\mathrm{i}\tau)\,G^{0}_{jk}(\mathbf{R}_{1}-\mathbf{R},-\mathrm{i}\tau)$
$\qquad+G^{0}_{ji}(-\mathbf{R},-\mathrm{i}\tau)\,G^{0}_{kl}(\mathbf{R}_{2}-\mathbf{R}_{1},\mathrm{i}\tau)$
$\qquad+G^{0}_{li}(-\mathbf{R}_{2},-\mathrm{i}\tau)\,G^{0}_{kj}(\mathbf{R}-\mathbf{R}_{1},\mathrm{i}\tau)\Big].$ (19)
@@
In the $GW$ approximation, the self-energy can be decomposed into a time-independent exchange part $\Sigma^{\rm x}$ and a time-dependent correlation part $\Sigma^{\rm c}$, i.e., $\Sigma=\Sigma^{\rm x}+\Sigma^{\rm c}$. These two parts correspond to splitting the screened Coulomb interaction $W$ into the bare Coulomb interaction $V$ and a remaining dynamic part $W^{c}$, i.e., $W=V+W^{c}$. Accordingly, we can compute these two parts within the NAO basis as follows:
$\Sigma_{pp^{\prime}}^{\mathrm{x}}(\mathbf{k})$ $=\sum_{ij}c_{p\mathbf{k}}^{i*}c_{p^{\prime}\mathbf{k}}^{j}\sum_{\mathbf{R}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{R}}\Sigma_{ij}^{\mathrm{x}}(\mathbf{R}),$ (22)
$\Sigma_{pp^{\prime}}^{\mathrm{c}}(\mathbf{k},\mathrm{i}\tau)$ $=\sum_{ij}c_{p\mathbf{k}}^{i*}c_{p^{\prime}\mathbf{k}}^{j}\sum_{\mathbf{R}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{R}}\Sigma_{ij}^{\mathrm{c}}(\mathbf{R},\mathrm{i}\tau).$
Efficient algorithms to calculate the exact-exchange matrix in real space, $\Sigma_{ij}^{\mathrm{x}}(\mathbf{R})$, have been extensively discussed previously Lin et al. (2020, 2021, 2025), and will not be further discussed here. As for the correlation part, the matrix elements $\Sigma^{\rm c}_{ij}(\mathbf{R},\mathrm{i}\tau)$ can be expressed in terms of the LRI coefficients, the screened Coulomb interaction $W_{\mu\nu}^{c}(\mathbf{R},\mathrm{i}\tau)$, and the Green’s function as:
@@ III.1 QSGW in LibRPA
Subsequently, the QSGWGW iteration proceeds analogously to a standard $G^{0}W^{0}$ calculation in LibRPA until the self-energy on the imaginary frequency axis, $\Sigma_{pp^{\prime}}(\mathbf{k},\mathrm{i}\omega)$, is obtained.
A Padé approximation is then employed to perform the analytic continuation from the imaginary frequency axis to the real frequency axis, yielding the self-energy at real frequencies, $\Sigma_{pp^{\prime}}(\mathbf{k},\omega)$. A detailed discussion on the stability of this analytic continuation is provided in Sec. III.2.
@@
To improve the stability of the self-consistency cycle for molecular systems, we employ Pulay’s DIIS mixing method Pulay (1980) when updating the QSGWGW Hamiltonian. For periodic systems, the iterations are generally stable in our benchmarks and converge without DIIS mixing. Our QSGWGW implementation typically requires about 10 iterations to achieve convergence of eigenvalues within $10^{-3}$ eV for periodic systems, whereas molecular systems may still require tens to hundreds of iterations. Upon convergence, the final quasiparticle energies and orbitals are utilized to compute various electronic properties, such as band structures and topological invariants.
@@ III.2 Analytic continuation of self-energy
Figure 2: Basis representation dependence tests of analytic continuation in QSGWGW for “Mode A” and “Mode B” for Si (top) and MgO (bottom). The Si calculations employed a $6\times 6\times 6$ k-point mesh, while MgO employed an $8\times 8\times 8$ mesh. Both used a frequency grid of 16 points and the intermediate_gw basis sets. In the legend, “initial” denotes analytic continuation performed in the initial KS-DFT eigenbasis, while “updated” denotes analytic continuation performed in the updated QSGWGW eigenbasis from the previous iteration.
@@ III.3.1 $\mathbf{k}$-point and frequency grid convergence
First, taking Si as an example, we check the convergence of the QSGWGW band gap with respect to the Brillouin-zone sampling using uniform $n\times n\times n$ $\bf k$}-point meshes (n=6n=6–10), and the results are shown in Fig. 3. Taking the $10\times 10\times 10$ mesh as the reference, the residual QSGWGW band-gap difference decreases from 32 meV at $6\times 6\times 6$ to 6.5 meV at $9\times 9\times 9$ (with $E_{g}^{\mathrm{QS}GW}=1.3621$ eV at $9\times 9\times 9$ and $E_{g}^{\mathrm{QS}GW}=1.3686$ eV at $10\times 10\times 10$). At $8\times 8\times 8$ $\bf k$}-mesh, the difference from the reference value at $10\times 10\times 10$ is 12.5 meV. This level of convergence is not fully satisfactory, but it represents a good compromise between numerical accuracy and computational load; hence $8\times 8\times 8$ $\bf k$}-mesh is adopted in this work for cubic systems in production calculations.
In the present implementation, a likely reason for the relatively slow convergence with respect to $\bf k$}-mesh is that the head correction term (i.e., the $\Gamma$-point correction to the dielectric matrix; see Ref. Zhang et al. (2026)) is updated only in the first QSGWGW iteration and is kept fixed in subsequent iterations. We are still working on this issue by updating the treatment of the head correction term at each iteration and expect that this will improve the behavior of the $\bf k$}-point convergence.
@@ III.3.2 Basis sets
Here, intermediate_gw and really_tight denote the predefined FHI-aims species_defaults setting (defaults_2020), which explicitly specify (i) the orbital NAO basis (minimal basis plus additional confined radial functions, organized in tiers) and (ii) the associated numerical settings (integration grids and Hartree-potential accuracy) for each chemical element. In the defaults_2020 hierarchy, intermediate setting retains most numerical settings from tight but employs a more economical orbital basis, whereas tight setting corresponds to tier-based NAO orbital lists (e.g., “tier 2” for H–Ne and “tier 1” for most heavier elements), and really_tight setting keeps the same orbital basis as tight setting but uses tightly converged numerical parameters (including basis_dep_cutoff=0).FHI-aims team (2025)
For periodic $GW$ calculations, the corresponding intermediate_gw setting is identical to the standard intermediate setting except that it augments the ABFs used in the LRI (RI-LVL) representation of the Coulomb operator by adding extra for_aux functions (specifically, for_aux hydro 4 f 0.0); these for_aux functions are used only to generate the auxiliary basis and do not enter the KS orbital basis.FHI-aims team (2025); Ren et al. (2021)
@@
MRE (%) +2.12
MARE (%) 3.64
Table 1: Experimental vertical ionization potentials (IP, in eV), together with our QSGWGW results and independent QSGWGW reference values from two GW100 external datasets (ADF STO extra and TM def2-TZVPP). For GWGW-based methods, IPs correspond to $\varepsilon_{\mathrm{H}}^{\mathrm{QP}}$, where $\varepsilon_{\mathrm{H}}^{\mathrm{QP}}$ is the converged HOMO quasiparticle energy. Experimental IPs are those compiled in Ref. van Setten et al. (2015) (Table 2 therein; values marked * are vertical, following that compilation). Each molecule employs the really_tight NAO basis tier and 32 frequency points. The external QSGWGW reference values are taken from the GW100 database 33 (see main text for details). Here MARE and MRE denote the mean absolute relative error and mean relative error (in %), respectively.
Both settings correspond to predefined FHI-aims NAO tiers and have been shown to yield well-converged $G^{0}W^{0}$ quasiparticle gaps for NAO-based all-electron implementations on similar benchmark sets (except for ZnO due to significant basis incompletion error, will be discussed below). We therefore adopt the same basis tiers for QSGWGW to provide a consistent and well-tested numerical baseline for the present benchmarks.
@@ IV.1 Molecular Benchmarks
To validate our QSGWGW implementation for molecular systems, we perform benchmark calculations on a representative selection of small molecules from the GW100 test set van Setten et al. (2015); Kaplan et al. (2016); Caruso et al. (2016); van Setten et al. (2013) (Table 1). The selected molecules cover diverse bonding characters and chemical environments, including a metallic dimer (Ag2), an ionic fluoride (AlF3), $p$-block elemental/covalent species (As2, AsH3), a polar diatomic (BF), an electron-deficient hydride (BH3), and a small $\pi$-bonded hydrocarbon (C2H2). Together, they provide a compact yet nontrivial validation set for the present implementation.
@@
In Table 1, the GW100 reference datasets are labeled by the underlying code and basis settings. In particular, “ADF STO extra” refers to the QSGWGW HOMO dataset computed with the Amsterdam Density Functional (ADF) code using a Slater-type orbital (STO) basis in the larger “extra” basis sets Förster and Visscher (2021) as provided in the GW100 database (GW100 file qsGW_HOMO_ADF_STO_extra.json). Likewise, “TM def2-TZVPP” refers to the QSGWGW HOMO dataset computed with TURBOMOLE (TM) using the Gaussian def2-TZVPP basis sets Caruso et al. (2016) as provided in the GW100 database (GW100 file qsGW_HOMO_Tv6.0_def2-TZVPP.json).
@@ IV.2 Semiconductors and Insulators
wAlN wurtzite 3.110 4.980 0.3820
wZnO wurtzite 3.250 5.207 0.3825
@@
Table 2: Crystal structure prototypes and lattice constants (in Å) used for the periodic benchmark set in Table 3. For cubic systems, $c$ is not applicable. For wurtzite systems, $u$ denotes the internal fractional $z$ coordinate of the anion on the Wyckoff site $(1/3,2/3,u)$ (with the second anion at $(2/3,1/3,u+1/2)$).
Table 3 presents the band gaps of a series of semiconductors and insulators computed by our NAO-based QSGWGW implementation. We first validate our approach against established implementations by comparing our QSGWGW gaps to literature benchmarks from independent codes (Table 3). In particular, we list QSGWGW reference values obtained with crystalline Gaussian basis sets Lei and Zhu (2022) and all-electron LAPW-based QSGWGW benchmark values Salas-Illanes et al. (2022). For additional context on the spread across widely used implementations, we also list representative self-consistent $GW$-family reference results obtained within a PMT (plane waves plus muffin-tin orbitals) hybrid-QSGWGW workflow Deguchi et al. (2016), a PAW-based fully self-consistent $GW$ implementation Grumet et al. (2018), and a plane-wave self-consistent $GW$ scheme with efficient vertex corrections Chen and Pasquarello (2015).
To facilitate cross-code validation, we adopt the Gaussian-QSGWGW Lei and Zhu (2022) and LAPW-QSGWGW Salas-Illanes et al. (2022) results listed in Table 3 as our primary QSGWGW reference benchmarks where available. The remaining entries—PMT hQSGWGW Deguchi et al. (2016), PAW scGWGW Grumet et al. (2018), and PW scGWGW+vertex Chen and Pasquarello (2015)—represent closely related self-consistent $GW$-type approaches. These are included to illustrate the spread among widely used methodologies. However, since they are not strictly equivalent to the standard QSGWGW formalism, they are treated as contextual references rather than direct benchmarks. The superscript $\mathrm{zb}$ indicates that the corresponding value for ZnO was reported for the zinc-blende structure in Ref. Grumet et al. (2018), as compiled in Ref. Lei and Zhu (2022).
@@
Gaussian QSGWGW Lei and Zhu (2022) LAPW QSGWGW Salas-Illanes et al. (2022) PMT hQSGWGW Deguchi et al. (2016) PAW scGWGW Grumet et al. (2018) PW scGWGW+vertex Chen and Pasquarello (2015) FP-LMTO QSGWGW Kotani et al. (2007)
@@
Table 3: Band gaps (in eV) of semiconductors and insulators. An $8\times 8\times 8$ k-point mesh was employed for most systems, with the exception of wAlN and wZnO ($8\times 8\times 5$). All calculations used a frequency grid of 16 points and the intermediate_gw basis sets (see Sec. III.3.2). Experimental band gaps are taken from Ref. Ren et al. (2021) and references therein, except for Ne, which is from Ref. Schwentner et al. (1975). For BP, reported experimental values range from 2.2 to 2.5 eV; we use the midpoint value of 2.35 eV when computing error metrics. Literature reference values are cited in the column headers, with further interpretation provided in the main text.
@@
As illustrated in Fig. 5, the PBE functional systematically underestimates the band gaps for this benchmark set. The single-shot $G^{0}W^{0}$@PBE method captures most of the gap increase relative to PBE, while QSGWGW provides a further systematic gap opening on top of $G^{0}W^{0}$@PBE. Under the present computational protocol, QSGWGW overestimates the band gaps by roughly the same amount as $G^{0}W^{0}$@PBE underestimates them, when compared to the experimental values. In the following, we validate our QSGWGW results against independent QSGWGW benchmarks from the literature, with experimental gaps included only for contextual reference.
@@
First, we assess cross-implementation consistency by comparing our QSGWGW band gaps with literature benchmarks from independent implementations, specifically Gaussian-QSGWGW Lei and Zhu (2022) and LAPW-QSGWGW Salas-Illanes et al. (2022) (see Table 3). Across the set of materials common to both studies, our QSGWGW results typically agree within $\sim$0.3 eV with these literature values. Especially for wide-gap insulators like C (diamond) and BN, our results (6.41 eV and 7.29 eV, respectively) are fairly close to the literature values listed in Table 3. For the rare gas solid Ne, our QSGWGW result (22.84 eV) is in excellent agreement with the periodic QSGWGW benchmark value (22.57 eV) reported in Ref. Lei and Zhu (2022) and with the experimental value (21.7 eV). Overall, the observed cross-code differences are comparable to the spread reflected in reported periodic QSGWGW results, supporting the validity of our NAO-based QSGWGW scheme relative to established implementations.
The largest deviations are observed for wurtzite ZnO (wZnO) (3.71 eV), which is 0.92 eV lower than the Gaussian-QSGWGW value (4.63 eV), and for MgO (9.14 eV), which is 0.83 eV lower than the LAPW-QSGWGW value (9.97 eV). wZnO is a system known to be challenging for the $GW$ methods, due to the strong hybridization between Zn $3d$ states and O $2p$ states, as well as the slow convergence with respect to high-energy unoccupied states Friedrich et al. (2011). Previous studies have shown that, in the LAPW framework, adding high-energy local orbitals (HLOs) can significantly increase the $GW$ band gap of ZnO Jiang and Blaha (2016). The underestimation of the $G^{0}W^{0}$ band gap for ZnO within the standard NAO basis sets has been discussed in Ref. Ren et al. (2021). It has been shown Ren et al. (2021) that the issue can be mitigated by complementing NAOs with highly localized Slater-type orbitals (STOs), in analogy with the LAPW+HLOs scenario.